# Learning-based agricultural management in partially observable environments subject to climate variability

### v1

@@ 1 Introduction
According to a 2022 report from the United States Department of Agriculture (USDA) Wang2022 , total farm production nearly tripled from 1948 to 2017. However, despite the growth, there remains a global food shortage. The Food and Agriculture Organization (FAO) estimated that approximately 828 million people were experiencing hunger in 2022. Given this pressing issue, it becomes imperative to leverage new technologies to boost farm production, and one such solution is Precision Agriculture (PA) Zhang2022 . Precision agriculture, also known as “precision farming” or “prescription farming,” utilizes information and technology-based agricultural management systems. These systems enable farmers to precisely tailor their soil and crop management practices to various weather/soil conditions on individual farmlands.
@@
In a modern community, PA is an emerging field aimed at enhancing the efficiency and sustainability of agricultural practices Duru2015 . Precision agriculture often employs advanced technologies such as remote sensing, robotics, Machine Learning (ML), and Artificial Intelligence (AI) techniques. Monitoring plant health and detecting diseases are vital aspects of sustainable agriculture. Yet, manual disease detection is labor-intensive, necessitating significant expertise, effort, and extended processing time. Researchers have turned to image recognition algorithms as a solution, achieving promising results in plant disease identification Khirade2015 . Additionally, AI’s potential in forecasting crop yields has gained significant attention. Some researchers have used satellite imagery to develop models that predict yields, often incorporating crop identification maps and meteorological data. These models have been applied to forecast yields for crops such as wheat, rice, cotton, and sugarcane, especially in regions like the Indus Basin in Pakistan, demonstrating satisfactory performance Bastiaanssen2003 .
@@
As one of the important components in PA, learning-based agricultural management represents a substantial departure from traditional farming methods, which often rely on human intuition and experience. Learning-based agricultural management adopts a more data-driven approach Kirtan2019 with the overarching goals of increasing efficiency, reducing waste, protecting the environment, and improving the sustainability of farming practices. A notable example is seen in the work of Vij and co-authors Vij2020 , who predicted the irrigation needs of farmland. They achieved this by using intelligent systems to monitor ground parameters, including soil moisture, soil temperature, and environmental conditions such as air temperature, ultraviolet rays, light radiation, and the relative humidity of the fields. Additionally, they incorporated weather forecast data sourced from the internet.
@@
In previous studies of agricultural management, researchers traditionally collected and analyzed historical data to identify empirical regularities, which could then inform future agricultural policies and practices Kenneth2002 . With the continuous advancement of computer simulation technology, specialized software tools such as Decision Support System for Agrotechnology Transfer (DSSAT) Jones2003 , Agricultural Production Systems Simulator (APSIM) Keating2003 , and AquaCrop Steduto2009 have been developed. These simulation tools are designed to model various aspects of crop growth, yield, water, and nutrient requirements in response to environmental conditions. Particularly, DSSAT has gained widespread recognition and has been employed for over 30 years to simulate crop behavior and responses to environmental variables, making it a valuable resource for crop simulation studies.
@@
The above-mentioned software tools can effectively approximate the growth process of crops and predict the final yields by considering management parameters and environmental conditions such as temperature, humidity, soil property, and other influential factors. Among those factors, nitrogen fertilizer stands out as a crucial element that can be managed and controlled. Nitrogen is the primary nutrient that profoundly affects crop growth and yield production. However, an excessive application of nitrogen fertilizer can lead to substantial detrimental effects on the environment Mark2011 , including nitrate leaching. Therefore, it becomes imperative to implement effective nitrogen management strategies to balance optimizing crop yields, minimizing environmental damage, and sustaining farmers’ income.
@@
As one subset of ML, Reinforcement Learning (RL) empowers computer programs, acting as agents, to control unknown and uncertain dynamical systems while pursuing specific tasks Li2023 ; Cai2023 . This approach has garnered increasing attention from researchers interested in determining optimal strategies for agricultural management Overweg2021 . Notably, the DSSAT, a widely recognized agricultural simulation tool Jones2003 , has been extended to a realistic simulation environment known as Gym-DSSAT Romain2022 . In this environment, RL agents can effectively learn fertilization and irrigation management strategies when provided with soil property data and weather history/forecast information. Specifically, Gautron et al. Romain2022 proved that RL agents could discover interesting crop management policies in simulated conditions and gym-DSSAT and simulated worldwide growing conditions. In addition, Wu et al. Wu2022 recently demonstrated that the RL-trained policies outperformed empirical methods, resulting in higher or similar crop yields while using fewer fertilizers. Sun et al. Sun2017 conducted research wherein they formulated a reinforcement learning-driven irrigation control method. This technique has the potential to substantially enhance net gains by accounting for both crop yield and water expenditure.
@@
The aforementioned works Romain2022 ; Wu2022 ; Sun2017 predominantly assumed the agricultural environment was fully observable, leading to the mathematical formulation of corresponding RL problems as Markov Decision Process (MDP) problems. In MDP, each state of the environment is expected to encompass all the necessary information for the agent to determine the best action for optimizing the objective function. However, questions arise regarding whether the state variable listed in Gym-DSSAT can comprehensively represent the state of the agricultural environment Puterman2014 . Moreover, certain state variables, such as the index of plant water stress, daily nitrogen denitrification, and daily nitrogen plant population uptake, may be challenging to measure and access.
@@
This issue mirrors many real-world applications where agents lack complete knowledge to determine the environment’s state precisely. In such cases, agents often only have access to uncertain or incomplete observations of the states. This challenge may be addressed by the Partially Observable Markov Decision Process (POMDP) framework Astrom1965 . While POMDP was mentioned in the context of Gym-DSSAT Romain2022 during its introduction, the specific solution was not detailed. Recently, Tao et al. Tao2022 employed Imitation Learning (IL) to develop management policies that require only a minimal number of state variables by mirroring the actions of the RL policies learned with full observation. They discovered that the policies, after being learned under partial observation, demonstrated decisions almost identical to those trained with RL under full observation.
@@
On the other hand, climate variability is another critical factor in agriculture and its management, encompassing changes in temperature, precipitation, wind patterns, and other meteorological elements occurring over various temporal and spatial scales IPCC2014 . Weather conditions, especially extreme weather events, can significantly impact final crop yields. For example, Motha and Baier Motha2005 conducted a study analyzing the time series of corn yields from 1895 to 2002 in the state of Iowa. They identified substantial agricultural losses in 1988 due to one of the worst droughts during the growing season in modern history. Additionally, flooding caused almost a 50% drop in Iowa’s corn production in 1993 compared to the previous year. Therefore, it is crucial for the learning agent of agricultural management to be adaptive to climate variability.
@@
First, our study demonstrates the effectiveness of formulating the agricultural environment as a POMDP in generating superior policies (i.e., management strategies) compared to using an MDP, which has been the assumption in most prior works Romain2022 ; Wu2022 ; Sun2017 . This conclusion contrasts with the findings of Tao et al. Tao2022 , where the agent was initially pre-trained in MDP and subsequently in POMDP through imitation learning, resulting in similar policy performance. Furthermore, our approach enables the agent to learn optimal policies within the POMDP framework directly. Specifically, we employ a model-free RL method that incorporates RNNs to solve POMDP problems.
@@ 2.1 MDP and POMDP
Definition 2.2 (POMDP).
A POMDP can be generally denoted by a tuple $\mathcal{P}=\left(S,A,T,s_{0},R,O,\Omega\right)$, where $S,A,T,S_{0},$ and $R$ are defined as the same in MDP (Definition 2.1), and
  • 1.
@@
There have been several model-based approaches Li2023 to solving POMDP problems. These approaches commonly seek optimal policies in the belief state space rather than the state space defined in the POMDP (Definition 2.2). A belief state is a probability distribution encompassing all the possible states where the agent could be. It can be dynamically updated based on transition and observation probabilities during the learning process. When using the model-based approach, a POMDP problem transforms into a search for optimal policies within a corresponding MDP defined in the belief state space. However, in this study, the agent lacks knowledge of the transition and observation probabilities, rendering model-free RL methods is an appropriate choice.
@@ 2.2 Q-learning and Deep Q-Network
Q-learning Watkins1992 is a model-free, value-based RL method in which the agent tends to achieve optimal state-action values. As a tabular method, the naïve Q-learning employs a Q-table to store $Q$ values and quantify the best action with the highest $Q$ value for the agent to choose. On the other hand, $Q$ values in the Q-table are updated via bootstrapping when the agent interacts with its environment. In each episode, $Q(s,a)$ is updated at each step as below after taking action $a$ at state $s$, following the Bellman equation Sutton2018 .
@@
However, tabular Q-learning becomes unsuitable when dealing with a large or infinite state space, such as agricultural environments. This challenge can be addressed by replacing the Q-table with deep neural networks (DNNs), known as Q-networks, to estimate $Q$ values. This approach falls under the umbrella of Deep Reinforcement Learning (DRL) Arulkumaran2017 , and in this study, we employed the Deep Q-Network (DQN) Mnih2013 , which is an extension of Q-learning.
Deep Q-Network consists of two Q-networks: an evaluation Q-network, denoted as $Q_{e}(s,a;\theta_{e})$, and a target Q-network, denoted as $Q_{t}(s,a;\theta_{t})$. Here, $\theta_{e}$ and $\theta_{t}$ represent the network weights, which are to be trained and updated through the experience replay memory Lin1992 . During the learning process, an experience is generated at each step in the form of $(s,a,s^{\prime},R)$ and stored in a memory pool. Simultaneously, a set of these experiences, referred to as a mini-batch, is selected from the memory pool to train and update the evaluation Q-network. The Bellman equation presented in Equation (2) is modified as follows.
@@
In MDP, the agent has a complete knowledge of the environment, and the value function assesses available actions at the current state for decision-making. However, when the environment becomes partially observable, the agent must rely on a history of perceived information, typically a sequence of observations, to make informed decisions. As a result, the Q-networks in DQN take this sequence of observations as the input and produce corresponding $Q$ values. Furthermore, the policy or agent function now maps a sequence of observations to an action. In such cases, RNNs emerge as strong candidates for Q-networks because of their proficiency in handling sequential or time-series data.
In a recent study by Li et al. Li2023 , an RNN-based DQN was proposed for robotics motion planning in partially observable environments. They utilized Long Short-Term Memory (LSTM) Hochreiter1997 in their Q-networks to process sequences of observations. In this study, we opt for Gated Recurrent Units (GRU) Cho2014 , another advanced RNN architecture, within our Q-networks to model temporal dependencies among observations, incorporating feedback loops within the network structure. We conduct a comparison between LSTM-based DQN and GRU-based DQN, and both approaches yield similar results. However, due to its forget gate and the absence of an output gate, a GRU cell has fewer parameters, which results in increased efficiency and reduced training time compared to LSTM.
@@ 3 Agriculture management as a POMDP problem
In this study, we utilize maize crop growth in Iowa as a case study to demonstrate that the agricultural environment, represented by DSSAT, is partially observable. Our study encompasses the years 1965, 1980, 1999, and 2020. We obtain the corresponding weather data from the Iowa State University Soil Moisture Network ISUSMN , which includes daily maximum temperature, minimum temperature, solar radiation, and precipitation. However, due to a lack of comprehensive data, we rely on the soil property data from 1999 and apply it to the other years in our study. The basic Gym-DSSAT input file from 1999 weather and soil data and the DRL code for this research can be found in our GitHub repository. (https://github.com/ZhaoanWang/Learning-based-agricultural-Management).
@@ 3.1 Model setup
The Gym-DSSAT employs factored representations, utilizing a total of 28 internal variables, as detailed in Table 1. Many studies Romain2022 ; Wu2022 have used these variables to represent the agricultural environment’s state. They have framed learning-based agricultural management problems as MDP problems, assuming full observability of the environment. This assumption implies that the environment possesses the Markov property, allowing the agent to make decisions based on the immediately-received state variables. However, it is important to note that there is no conclusive evidence to demonstrate that these 28 internal variables can entirely determine the state of the agricultural environment. Furthermore, not all of these variables are easily observable or accessible. This study uses the first 10 variables from Table 1 as observation variables. This approach enables the agent to make decisions based on both current and previous observations. Consequently, the original problem is transformed into a POMDP problem.
Variable
Description
cumsumfert
cumulative nitrogen fertilizer applications (kg/ha)
dap
days after planting
istage
DSSAT maize growing stage
pltpop
[…]
rain
rainfall for the current day (mm/d)
sw
[…]
tmax
maximum temperature for the current day (${}^{\circ}$)
tmin
minimum temperature for the current day (${}^{\circ}$)
vstage
vegetative growth stage (number of leaves)
xlai
plant population leaf area index
cleach
cumulative nitrate leaching (kg/ha)
cnox
cumulative nitrogen denitrification (kg/ha)
dtt
growing degree days for the current day (C/d)
es
actual soil evaporation rate (mm/d)
grnwt
grain weight dry matter (kg/ha)
nstres
index of plant nitrogen stress
pcngrn
massic fraction of nitrogen in grains
rtdep
root depth (cm)
runoff
calculated runoff (mm/d)
srad
[…]
swfac
index of plant water stress
tleachd
daily nitrate leaching (kg/ha)
tnoxd
daily nitrogen denitrification (kg/ha)
topwt
above the ground population biomass (kg/ha)
totir
total irrigated water (mm)
trun
daily nitrogen plant population uptake (kg/ha)
wtdep
depth to water table (cm)
wtnup
cumulative plant population nitrogen uptake (kg/ha)
@@
This section explores four problems using different MDP and POMDP models. We then compare corn yields and N fertilizer usages resulting from the optimal policies {-}- i.e., management strategies learned by the agent {-}- in the previously mentioned years. These problem types are as follows:
  • 1.
@@
Given that maize crops in Iowa are typically rain-fed Wu2022 , this study excludes daily irrigation considerations and concentrates on nitrogen fertilization. Consequently, the action space encompasses various quantities of nitrogen that can be applied in a single day. Mathematically, the action space is discretized as ${10k(kg/ha)}$ nitrogen input, where $k$ ranges from 1 to 20.
In a given day ‘$d_{t}$,’ after taking action, which involves applying an amount of nitrogen $N_{t}$, the agent receives a reward defined as:
@@
where $Y$ represents the corn yield at harvest, and $L_{t}$ denotes nitrate leaching on a particular day $t$. The weight coefficients, $w_{1}$ and $w_{2}$, are determined by the prevailing prices of corn and nitrogen input in each simulated year, as listed in Table 2. Particularly, $w_{1}$ corresponds to the price of corn per kilogram USDA , and $w_{2}$ is based on the price of nitrogen per kilogram, calculated using 45% urea nitrogen to determine the price of 100% nitrogen ISU . In addition, $w_{3}$ is the weight assigned to nitrate leaching, calculated as a multiple of $w_{2}$. The specific multiple is 5, as indicated in Wu2022 .
Year
w1
w2
w3
1965
0.03819
0.26
1.04
1980
0.07953
0.49
1.96
1999
0.07087
0.39
1.95
2020
0.1827
0.87
3.48
@@
The training process includes 6000 episodes, each lasting 180 steps (i.e., days). We set the discount factor to 0.99. To update the neural networks, we utilize Pytorch and Adam optimizer Kingma2014 with an initial learning rate of 1e-5 and a batch size of 640. The simulations are conducted on two machines: one equipped with an Intel Core i7-12700K processor, NVIDIA GeForce RTX 3070 Ti graphics card, and 64GB RAM, while the other featured an AMD 5800h processor, NVIDIA GeForce RTX 3070 graphics card, and 32GB RAM.

### v2

@@ 1 Introduction
According to a 2022 report from the United States Department of Agriculture (USDA) [1], total farm production nearly tripled from 1948 to 2017. However, despite the growth, there remains a global food shortage. The Food and Agriculture Organization (FAO) estimated that approximately 828 million people were experiencing hunger in 2022. Given this pressing issue, it becomes imperative to leverage new technologies to boost farm production, and one such solution is Precision Agriculture (PA) [2]. Precision agriculture, also known as “precision farming” or “prescription farming,” utilizes information and technology-based agricultural management systems. These systems enable farmers to precisely tailor their soil and crop management practices to various weather/soil conditions on individual farmlands.
@@
In a modern community, PA is an emerging field aimed at enhancing the efficiency and sustainability of agricultural practices [3]. Precision agriculture often employs advanced technologies such as remote sensing, robotics, Machine Learning (ML), and Artificial Intelligence (AI) techniques. Monitoring plant health and detecting diseases are vital aspects of sustainable agriculture. Yet, manual disease detection is labor-intensive, necessitating significant expertise, effort, and extended processing time. Researchers have turned to image recognition algorithms as a solution, achieving promising results in plant disease identification [4]. Additionally, AI’s potential in forecasting crop yields has gained significant attention. Some researchers have used satellite imagery to develop models that predict yields, often incorporating crop identification maps and meteorological data. These models have been applied to forecast yields for crops such as wheat, rice, cotton, and sugarcane, especially in regions like the Indus Basin in Pakistan, demonstrating satisfactory performance [5].
@@
As one of the important components in PA, learning-based agricultural management represents a substantial departure from traditional farming methods, which often rely on human intuition and experience. Learning-based agricultural management adopts a more data-driven approach [6] with the overarching goals of increasing efficiency, reducing waste, protecting the environment, and improving the sustainability of farming practices. A notable example is seen in the work of Vij and co-authors [7], who predicted the irrigation needs of farmland. They achieved this by using intelligent systems to monitor ground parameters, including soil moisture, soil temperature, and environmental conditions such as air temperature, ultraviolet rays, light radiation, and the relative humidity of the fields. Additionally, they incorporated weather forecast data sourced from the internet.
@@
In previous studies of agricultural management, researchers traditionally collected and analyzed historical data to identify empirical regularities, which could then inform future agricultural policies and practices [8]. With the continuous advancement of computer simulation technology, specialized software tools such as Decision Support System for Agrotechnology Transfer (DSSAT) [9], Agricultural Production Systems Simulator (APSIM) [10], and AquaCrop [11] have been developed. These simulation tools are designed to model various aspects of crop growth, yield, water, and nutrient requirements in response to environmental conditions. Particularly, DSSAT has gained widespread recognition and has been employed for over 30 years to simulate crop behavior and responses to environmental variables, making it a valuable resource for crop simulation studies.
@@
The above-mentioned software tools can effectively approximate the growth process of crops and predict the final yields by considering management parameters and environmental conditions such as temperature, humidity, soil properties, and other influential factors. Among those factors, nitrogen fertilizer stands out as a crucial element that can be managed and controlled. Nitrogen is the primary nutrient that profoundly affects crop growth and yield production. However, an excessive application of nitrogen fertilizer can lead to substantial detrimental effects on the environment [12], including nitrate leaching. Therefore, it becomes imperative to implement effective nitrogen management strategies to balance optimizing crop yields, minimizing environmental damage, and sustaining farmers’ income.
@@
As one subset of ML, Reinforcement Learning (RL) empowers computer programs, acting as agents, to control unknown and uncertain dynamical systems while pursuing specific tasks [13, 14]. This approach has garnered increasing attention from researchers interested in determining optimal strategies for agricultural management [15]. Notably, the DSSAT, a widely recognized agricultural simulation tool [9], has been extended to a realistic simulation environment known as Gym-DSSAT [16]. In this environment, RL agents can effectively learn fertilization and irrigation management strategies when provided with soil property data and weather history/forecast information. Specifically, Gautron et al. [16] proved that RL agents could discover interesting crop management policies in simulated conditions and gym-DSSAT and simulated worldwide growing conditions. In addition, Wu et al. [17] recently demonstrated that the RL-trained policies outperformed empirical methods, resulting in higher or similar crop yields while using fewer fertilizers. Sun et al. [18] conducted research wherein they formulated a reinforcement learning-driven irrigation control method. This technique has the potential to substantially enhance net gains by accounting for both crop yield and water expenditure.
@@
The aforementioned works [16, 17, 18] predominantly assumed the agricultural environment was fully observable, leading to the mathematical formulation of corresponding RL problems as Markov Decision Process (MDP) problems. In MDP, each state of the environment is expected to encompass all the necessary information for the agent to determine the best action for optimizing the objective function. However, questions arise regarding whether the state variable listed in Gym-DSSAT can comprehensively represent the state of the agricultural environment [19]. Moreover, certain state variables, such as the index of plant water stress, daily nitrogen denitrification, and daily nitrogen plant population uptake, may be challenging to measure and access.
@@
This issue mirrors many real-world applications where agents lack complete knowledge to determine the environment’s state precisely. In such cases, agents often only have access to uncertain or incomplete observations of the states. This challenge may be addressed by the Partially Observable Markov Decision Process (POMDP) framework [20]. While POMDP was mentioned in the context of Gym-DSSAT [16] during its introduction, the specific solution was not detailed. Recently, Tao et al. [21] employed Imitation Learning (IL) to develop management policies that require only a minimal number of state variables by mirroring the actions of the RL policies learned with full observation. They discovered that the policies, after being learned under partial observation, demonstrated decisions almost identical to those trained with RL under full observation.
@@
On the other hand, climate variability is another critical factor in agriculture and its management, encompassing changes in temperature, precipitation, wind patterns, and other meteorological elements occurring over various temporal and spatial scales [22]. Weather conditions, especially extreme weather events, can significantly impact final crop yields. For example, Motha and Baier [23] conducted a study analyzing the time series of corn yields from 1895 to 2002 in the state of Iowa. They identified substantial agricultural losses in 1988 due to one of the worst droughts during the growing season in modern history. Additionally, flooding caused almost a 50% drop in Iowa’s corn production in 1993 compared to the previous year. Therefore, it is crucial for the learning agent of agricultural management to be adaptive to climate variability.
@@
First, our study demonstrates the effectiveness of formulating the agricultural environment as a POMDP in generating superior policies (i.e., management strategies) compared to using an MDP, which has been the assumption in most prior works [16, 17, 18]. This conclusion contrasts with the findings of Tao et al. [21], where the agent was initially pre-trained in MDP and subsequently in POMDP through imitation learning, resulting in similar policy performance. Furthermore, our approach enables the agent to learn optimal policies within the POMDP framework directly. Specifically, we employ a model-free RL method that incorporates RNNs to solve POMDP problems.
@@ 2.1 MDP and POMDP
Definition 2 (POMDP).
A POMDP can be generally denoted by a tuple $\mathcal{P}=\left(S,A,T,s_{0},R,O,\Omega\right)$, where $S,A,T,S_{0},$ and $R$ are defined as the same in MDP (Definition 1), and
@@
There have been several model-based approaches [13] to solving POMDP problems. These approaches commonly seek optimal policies in the belief state space rather than the state space defined in the POMDP (Definition 2). A belief state is a probability distribution encompassing all the possible states where the agent could be. It can be dynamically updated based on transition and observation probabilities during the learning process. When using the model-based approach, a POMDP problem transforms into a search for optimal policies within a corresponding MDP defined in the belief state space. However, in this study, the agent lacks knowledge of the transition and observation probabilities, rendering model-free RL methods an appropriate choice.
@@ 2.2 Q-learning and Deep Q-Network
Q-learning [28] is a model-free, value-based RL method in which the agent tends to achieve optimal state-action values. As a tabular method, the naïve Q-learning employs a Q-table to store $Q$ values and quantify the best action with the highest $Q$ value for the agent to choose. On the other hand, $Q$ values in the Q-table are updated via bootstrapping when the agent interacts with its environment. In each episode, $Q(s,a)$ is updated at each step as below after taking action $a$ at state $s$, following the Bellman equation [24].
@@
However, tabular Q-learning becomes unsuitable when dealing with a large or infinite state space, such as agricultural environments. This challenge can be addressed by replacing the Q-table with deep neural networks (DNNs), known as Q-networks, to estimate $Q$ values. This approach falls under the umbrella of Deep Reinforcement Learning (DRL) [25], and in this study, we employed the Deep Q-Network (DQN) [26], which is an extension of Q-learning.
Deep Q-Network consists of two Q-networks: an evaluation Q-network, denoted as $Q_{e}(s,a;\theta_{e})$, and a target Q-network, denoted as $Q_{t}(s,a;\theta_{t})$. Here, $\theta_{e}$ and $\theta_{t}$ represent the network weights, which are to be trained and updated through the experience replay memory [27]. During the learning process, an experience is generated at each step in the form of $(s,a,s^{\prime},R)$ and stored in a memory pool. Simultaneously, a set of these experiences, referred to as a mini-batch, is selected from the memory pool to train and update the evaluation Q-network. The Bellman equation presented in Equation (2) is modified as follows.
@@
In MDP, the agent has complete knowledge of the environment, and the value function assesses available actions at the current state for decision-making. However, when the environment becomes partially observable, the agent must rely on a history of perceived information, typically a sequence of observations, to make informed decisions. As a result, the Q-networks in DQN take this sequence of observations as the input and produce corresponding $Q$ values. Furthermore, the policy or agent function now maps a sequence of observations to an action. In such cases, RNNs emerge as strong candidates for Q-networks because of their proficiency in handling sequential or time-series data.
In a recent study by Li et al. [13], an RNN-based DQN was proposed for robotics motion planning in partially observable environments. They utilized Long Short-Term Memory (LSTM) [29] in their Q-networks to process sequences of observations. In this study, we opt for Gated Recurrent Units (GRU) [30], another advanced RNN architecture, within our Q-networks to model temporal dependencies among observations, incorporating feedback loops within the network structure. We conduct a comparison between LSTM-based DQN and GRU-based DQN, and both approaches yield similar results. However, due to its forget gate and the absence of an output gate, a GRU cell has fewer parameters, which results in increased efficiency and reduced training time compared to LSTM.
@@ 3 Agriculture management as a POMDP problem
In this study, we utilize maize crop growth in Iowa as a case study to demonstrate that the agricultural environment, represented by DSSAT, is partially observable. Our study encompasses the years 1965, 1980, 1999, and 2020. We obtain the corresponding weather data from the Iowa State University Soil Moisture Network [31], which includes daily maximum temperature, minimum temperature, solar radiation, and precipitation. However, due to a lack of comprehensive data, we rely on the soil property data from 1999 and apply it to the other years in our study. The basic Gym-DSSAT input file from 1999 weather and soil data and the DRL code for this research can be found in our GitHub repository. (https://github.com/ZhaoanWang/Learning-based-agricultural-Management).
@@ 3.1 Model setup
The Gym-DSSAT employs factored representations, utilizing a total of 28 internal variables, as detailed in Table 1. Many studies [16, 17] have used these variables to represent the agricultural environment’s state. They have framed learning-based agricultural management problems as MDP problems, assuming full observability of the environment. This assumption implies that the environment possesses the Markov property, allowing the agent to make decisions based on the immediately-received state variables. However, it is important to note that there is no conclusive evidence to demonstrate that these 28 internal variables can entirely determine the state of the agricultural environment. Furthermore, not all of these variables are easily observable or accessible. This study uses the first 10 variables from Table 1 as observation variables. This approach enables the agent to make decisions based on both current and previous observations. Consequently, the original problem is transformed into a POMDP problem.
Variable Description
cumsumfert cumulative nitrogen fertilizer applications (kg/ha)
dap days after planting
istage DSSAT maize growing stage
pltpop plant population density (plant/m2)
rain rainfall for the current day (mm/d)
sw volumetric soil water content in soil layers (cm3 [water] / cm3 [soil])
tmax maximum temperature for the current day (C)
tmin minimum temperature for the current day (C)
vstage vegetative growth stage (number of leaves)
xlai plant population leaf area index
cleach cumulative nitrate leaching (kg/ha)
cnox cumulative nitrogen denitrification (kg/ha)
dtt growing degree days for the current day (C/d)
es actual soil evaporation rate (mm/d)
grnwt grain weight dry matter (kg/ha)
nstres index of plant nitrogen stress
pcngrn massic fraction of nitrogen in grains
rtdep root depth (cm)
runoff calculated runoff (mm/d)
srad solar radiation during the current day (MJ/m2/d)
swfac index of plant water stress
tleachd daily nitrate leaching (kg/ha)
tnoxd daily nitrogen denitrification (kg/ha)
topwt above the ground population biomass (kg/ha)
totir total irrigated water (mm)
trun daily nitrogen plant population uptake (kg/ha)
wtdep depth to water table (cm)
wtnup cumulative plant population nitrogen uptake (kg/ha)
@@
This section explores four problems using different MDP and POMDP models. We then compare corn yields and N fertilizer usages resulting from the optimal policies {-} i.e., management strategies learned by the agent {-} in the previously mentioned years. These problem types are as follows:
@@
Given that maize crops in Iowa are typically rain-fed [17], this study excludes daily irrigation considerations and concentrates on nitrogen fertilization. Consequently, the action space encompasses various quantities of nitrogen that can be applied in a single day. Mathematically, the action space is discretized as 10k(kg/ha){10k(kg/ha)} nitrogen input, where $k$ ranges from 1 to 20.
In a given day $d_{t}$,’ after taking action, which involves applying an amount of nitrogen $N_{t}$, the agent receives a reward defined as:
@@
where $Y$ represents the corn yield at harvest, and $L_{t}$ denotes nitrate leaching on a particular day $t$. The weight coefficients, $w_{1}$ and $w_{2}$, are determined by the prevailing prices of corn and nitrogen input in each simulated year, as listed in Table 2. Particularly, $w_{1}$ corresponds to the price of corn per kilogram [32], and $w_{2}$ is based on the price of nitrogen per kilogram, calculated using 45% urea nitrogen to determine the price of 100% nitrogen [33]. In addition, $w_{3}$ is the weight assigned to nitrate leaching, calculated as a multiple of $w_{2}$. The specific multiple is 5, as indicated in [17].
Year w1 w2 w3
1965 0.03819 0.26 1.04
1980 0.07953 0.49 1.96
1999 0.07087 0.39 1.95
2020 0.1827 0.87 3.48
@@
The training process includes 6000 episodes, each lasting 180 steps (i.e., days). We set the discount factor to 0.99. To update the neural networks, we utilize Pytorch and Adam optimizer [34] with an initial learning rate of 1e-5 and a batch size of 640. The simulations are conducted on a machine equipped with an Intel Core i7-12700K processor, an NVIDIA GeForce RTX 3070 Ti graphics card, and 64 GB of RAM.